Given the functions f(x)= x^2 - 6x - 4 and g(x)= x - 4, find f(g(4)) .

Answer by edjones(8007)

f(x)= x^2 - 6x - 4 and g(x)= x - 4
g(4)=4-4=0
f(0)=0-0-4=-4
Ed